# Adaptive Multi-Scale Integration Unlocks Robust Cell Annotation in Histopathology Images

### v1

@@ Abstract
Identifying cell types and subtypes from routine histopathology images is essential for advancing the computational understanding of human disease. Existing tile-based models can capture detailed nuclear morphology yet often fail to incorporate the broader tissue context that influences a cell’s function and identity. In addition, available human annotations are typically coarse-grained and unevenly distributed across studies, making fine-grained, subtype-level supervision difficult to obtain.
To address these limitations, we introduce NuClass, a pathologist workflow–inspired framework for cell-wise multi-scale integration of nuclear morphology and microenvironmental context. NuClass consists of two main components: Path local, which focuses on nuclear morphology from $224^{2}$-pixel crops, and Path global, which models the surrounding $1024^{2}$-pixel neighborhood. A learnable gating module adaptively balances local detail and contextual cues. To encourage complementary learning, we incorporate an uncertainty-guided objective that directs the global path to prioritize regions where the local path is uncertain, and we further provide calibrated confidence estimates and Grad-CAM visualizations to enhance interpretability.
To overcome the lack of high-quality annotations, we construct a marker-guided dataset from Xenium spatial transcriptomics assays, yielding single-cell–resolution labels for more than two million cells across eight organs and 16 classes. Evaluated on three fully held-out cohorts, NuClass achieves up to 96% F1 for its best-performing class, outperforming strong baselines. Our results demonstrate that multi-scale, uncertainty-aware fusion can bridge the gap between slide-level pathological foundation models and reliable, cell-level phenotype prediction.

### v2

@@ Abstract
Identifying cell types and subtypes in routine histopathology is fundamental for understanding disease. Existing tile-based models captures nuclear detail but miss the broader tissue context that influences cell identity. Current human annotations are coarse-grained and uneven across studies, making fine-grained, subtype-level classificsation difficult. In this study, we build a marker-guided dataset from Xenium spatial transcriptomics with single-cell–resolution labels for more than two million cells across eight organs and 16 classes to overcome the lack of high-quality annotations. Leveraging this data resource, we introduce NuClass, a pathologist workflow–inspired framework for cell-wise multi-scale integration of nuclear morphology and microenvironmental context. It combines Path local, which focuses on nuclear morphology from $224^{2}$-pixel crops, and Path global, which models the surrounding $1024^{2}$-pixel neighborhood, through a learnable gating module to balance local and global details. An uncertainty-guided objective directs the global path to prioritize regions where the local path is uncertain, and we provide calibrated confidence estimates and Grad-CAM maps for interpretability. Evaluated on three fully held-out cohorts, NuClass achieves up to 96% F1 for its best-performing class, outperforming strong baselines. Our results demonstrate that multi-scale, uncertainty-aware fusion can bridge the gap between slide-level pathological foundation models and reliable, cell-level phenotype prediction.